### v1

# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

Abstract
Satellites are increasingly adopting on-board Artificial Intelligence (AI) techniques to enhance platforms’ autonomy through edge inference. In this context, the utilization of deep learning (DL) techniques for segmentation in HS satellite imagery offers advantages for remote sensing applications, and therefore, we train 16 different models, whose codes are made available through our study, which we consider to be relevant for on-board multi-class segmentation of HS imagery, focusing on classifying oceanic (sea), terrestrial (land), and cloud formations. We employ the HYPSO-1 mission as an illustrative case for sea-land-cloud segmentation, and to demonstrate the utility of the segments, we introduce a novel sea-land-cloud ranking application scenario. Our system prioritizes HS image downlink based on sea, land, and cloud coverage levels from the segmented images. We comparatively evaluate the models for in-orbit deployment, considering performance, parameter count, and inference time. The models include both shallow and deep models, and after we propose four new DL models, we demonstrate that segmenting single spectral signatures (1D) outperforms 3D data processing comprising both spectral (1D) and spatial (2D) contexts. We conclude that our lightweight DL model, called 1D-Justo-LiuNet, consistently surpasses state-of-the-art models for sea-land-cloud segmentation, such as U-Net and its variations, in terms of performance (0.93 accuracy) and parameter count (4,563). However, the 1D models present longer inference time (15s) in the tested processing architecture, which is clearly suboptimal. Finally, after demonstrating that in-orbit image segmentation should occur post L1b radiance calibration rather than on raw data, we additionally show that reducing spectral channels down to 3 lowers models’ parameters and inference time, at the cost of weaker segmentation performance.
keywords:
Remote Sensing , Satellite Imagery , Hyperspectral Imaging , Deep Learning , Segmentation
[…]
[inst1]organization=Dept. Engineering Cybernetics, Norwegian University of Science and Technology,city=Trondheim, country=Norway,
[…]
[inst2]organization=Dept. of Computer Science, University of Bucharest,country=Romania,
[…]
[inst3]organization=Independent Consultant,city=Montreal, country=Canada,
1 Introduction
Artificial Intelligence (AI) has seen a growing utilization aboard satellite platforms for remote sensing, benefiting missions equipped with various technologies from conventional RGB to both Multispectral Imaging (MSI) and Hyperspectral Imaging (HSI). The importance of monitoring the Earth’s environment with HSI technology is emphasized by the use of hyperspectral (HS) imagers in numerous contemporary satellite missions, e.g., ESA/Phi-Sat-1 [1], ASI/PRISMA [2, 3], DLR/EnMAP [4], and NTNU/HYPSO-1 [5]. The importance of HS imagers on satellites is further highlighted by upcoming missions, e.g., NASA/PACE, GLIMR and SBG [6], ESA/CHIME [7], ASI/PRISMA Second Gen [8], as well as NTNU/HYPSO-2 and NTNU/HYPSO-3 [9]. HS missions offer great potential due to their detailed spectral characterization of Earth’s features, but they also share a common limitation with other optical remote sensing technologies: the challenge of cloud formation in the atmosphere obstructing the satellites’ capacity to monitor targets of interest. Techniques such as semantic segmentation proves valuable in a broad range of space applications, e.g., from cloud recognition to wildfire detection [10, 11, 12, 13, 14, 15]. Segmentation can indeed serve as a cloud detection technique, categorizing image pixels as either cloudy or non-cloudy, essentially performing binary classification. As a mission example, after the launch in 2020 of the Phi-Sat-1 HS mission part of ESA’s EOP initiative [1], a significant milestone was marked in 2021 when, for the first time, on-board AI was introduced in a satellite for object detection, with the mission’s initial focus being the detection of clouds aboard the satellite. Phi-Sat-1 marked the first in-orbit inference with its convolutional CloudScout Deep Learning (DL) model for cloud detection [10, 13]. The mission’s imager, HyperScout-2, captures the Visible and Near-Infrared (VNIR) and Thermal Infrared (TIR) spectral ranges, and subsequently deploys data processing via AI models in its on-board configurable platform facilitating updates for its algorithms in flight [1, 11]. These in-orbit updates rely on an embedded configurable Commercial off-the-shelf (COTS) platform, which is based on a Vision Processing Unit (VPU) [11, 16]. Furthermore, beyond the Phi-Sat-1 mission, the literature [12] also extends cloud segmentation to images of the missions Sentinel-2B (MSI) and FACSAT-1 (RGB) [17]. For instance, the state of the art [12] deploys a satellite data priority system to compute the cloud coverage level via DL-based segmentation on microcontrollers (MCUs), which determines the image downlink priority. Furthermore, the literature demonstrates that cloud detection can enhance data management aboard satellite platforms as presented next. On the one hand, selective HSI compression methods may be applied to cloud segments via CCSDS-123.0-B-2 to reduce their encoding rate [18, 19], while methods like Compressed Sensing (CS) eliminate on-board compression of cloudy data, shifting computation to ground-based reconstruction [20, 21]. In summary, segmentation can lead to selective compression, reducing the bit rate of data segregated as cloudy or its absolute removal by data screening [11]. Consequently, this streamlines postprocessing and enables the satellite to downlink only valuable non-cloudy data, thus optimizing the efficient usage of the usually limited communication channel [12]. On the other hand, as AI-driven cloud detection segregates non-cloud data from cloudy data, on-board cloud segmentation can additionally play a role in reducing decision-making latency, increasing satellite autonomy with more optimal on-board operations [11, 12, 14, 22, 23, 24, 25]. Furthermore, from a broader perspective, the same models initially trained for cloud binary segmentation can be retrained to solve more complex problems, e.g., multi-class segmentation, where each image pixel is assigned to one of several categories. To further illustrate the importance of AI-powered cloud and multi-class segmentation for many applications, we next employ a particular small satellite mission as an example case. Namely, for space initiatives such as HYPSO-1 (HYPerspectral small Satellite for Ocean observation) [5], the reduction in latency is especially important for time-sensitive remote sensing applications, such as early event warnings. HYPSO-1 focuses on detecting marine phenomena such as Harmful Algal Blooms (HABs), which can lead to significant fish farm salmon mortality. To this extent, segmentation for adequate on-board data management can play a crucial role in achieving prompt management responses to mitigate adverse effects in such scenarios. In the context of multi-class segmentation, by segregating the data along various categories of interest, on-board AI can properly enable the activation of other agents in observational hierarchies monitoring the environment at multiple spatial and temporal scales with drones and marine vessels [5, 26, 27], where the HYPSO-1 mission serves as the space segment of the observational pyramid. However, handling increasingly advanced tasks on satellites, such as segmentation of growing complexity, is feasible only if COTS are conveniently optimized. These hardware solutions provide benefits by accelerating, with high power efficiency, the repetitive operations commonly found in DL networks [1, 13, 16, 22]. Indeed, to enable AI applications in satellites, the literature explores various COTS hardware from the conventional MCUs [12] to the high-performance VPUs aboard Phi-Sat-1, Field-Programmable Gate Arrays (FPGAs) on board HYPSO-1, Tensor Processing Units (TPUs), Graphics Processing Units (GPUs), and other dedicated accelerators [1, 9, 10, 28, 29, 30].
To our knowledge, the existing literature in the field of cloud detection on board satellite platforms does not compare how different shallow and deep learning models, including distinctions like 1D vs. 2D-CNN-based models, influence the performance and complexity of cloud segmentation. Furthermore, the field also overlooks the effects of reducing the number of spectral channels, particularly in MSI/HSI data, on models for cloud segmentation. The existing literature lacks comparisons of key aspects, such as whether on-board AI for cloud detection, is better applied to raw data or L1b radiance data. It does not discuss how calibration impacts segmentation results and does not provide justification for the optimal stage in the on-board data processing pipeline for image segmentation. Moreover, within the literature pertaining to cloud detection in HS satellite imagery, the primary emphasis naturally lies on binary segmentation, specifically distinguishing between the presence of clouds and non-cloudy data; in fact, as earlier mentioned, there are ranking systems that assess cloud levels based on the segmented outcomes. Nevertheless, none of the current systems perform multi-class semantic segmentation for the on-board detection of clouds, sea, and land categories, which typically account for most of the variance in HS satellite imagery. The absence of sea-land-cloud segmentation systems impedes, as a result, the development of applications like priority ranking systems that can integrate these three categories into a unified ranker. In this work, we tackle all the mentioned research gaps and utilize the HYPSO-1 mission as a case study, using “The HYPSO-1 Sea-Land-Cloud-Labeled Dataset” openly accessible in our prior article [31]. Following a methodology inspired by the data science CRISP-DM approach (CRoss Industry Standard Process for Data Mining) to ensure robust and reliable results, we train and compare 16 different Machine Learning (ML) and DL models. This extends the initial model baseline for HYPSO-1 established in [31], and our research findings make contributions to the field of on-board AI in the following manners:
  • 1.
    We propose and fine-tune four new 1D-CNN and 2D-CNN models, with the aim of adapting them for on-board deployment on satellites. This optimization entails reducing the number of model parameters while maintaining performance.
  • 2.
    We demonstrate that 1D-CNN models exhibit superior performance when compared to 2D-CNNs. Notably, our optimized 1D-CNN models like 1D-Justo-LiuNet and 1D-Justo-HuNet consistently outperform other state-of-the-art models, particularly in the case context of multi-class sea-land-cloud HS segmentation for the HYPSO-1 mission.
  • 3.
    We substantiate that segmenting L1b radiance data, as opposed to using raw data, enhances the segmentation of the top-performing models.
  • 4.
    We explore the impact of channel reduction on sea-land-cloud segmentation models.
  • 5.
    Finally, we propose an image ranking system designed for downlink prioritization based on sea, land, and cloud coverage.
2 Methodological Background
2.1 Data Source: A Hyperspectral Labeled Datastet from HYPSO-1
We employ “The Sea-Land-Cloud-Labeled Dataset” [31], publicly accessible at https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset. The dataset comprises pixel-level labels for 25 million pixels across 38 scenes captured by HYPSO-1, where each HS image has spatial dimension of 956 x 684 pixels. Each pixel, with ground resolution of approximately 100 m x 600 m in the across- and along-track directions [5, 26], contains a spectral signature of 120 channels covering the 400 to 800 nm range in the VNIR spectrum, with 5 nm spectral resolution. The dataset consists of diverse scenes from various environments and provides raw data and L1b calibrated radiance for comparison with other HS satellites. As for the ground-truth labels, the dataset includes expert pixel-level annotations for three classes: “sea”, “land”, and “clouds”, which are basic categories due to the lack of on-site validation impeding the reliable labeling of more detailed oceanic and terrestrial subcategories. The dataset exhibits a slight class imbalance with 37.01% pixels representing sea, 40.14% representing land, and 22.85% representing clouds/overexposed areas.
2.2 Data Preprocessing: Dimensionality Reduction and Normalization
While our primary focus is HSI segmentation, rather than dimensionality reduction extensively studied for HSI [32], we still conduct spectral channel reduction as a data preprocessing step prior to training ML and DL models. This channel reduction is crucial when dealing with high-dimensional HSI data, as not all channels may contain valuable information. The method in this section helps identify bands for potential removal from the sea-land-cloud dataset. After dimensionality reduction, we apply min-max normalization to maintain a consistent data range and address any potential outlier effects.
Our channel reduction process begins with calculating the statistical standard deviation for each channel (L1b radiance), which provides an initial insight into potential unusual channels such as those with exceptionally low standard deviations demanding further investigation. In particular, we follow two approaches for detecting the most divergent points within the set of deviations shown in Fig. 1: visual inspection for potential outliers (e.g. extremely low deviation values), and the use of algorithms like the Isolation Forest (IF), which consists of an anomaly detection technique to assist identifying anomalous deviation points.
The IF algorithm employs a sensitivity threshold empirically set at 0.08 to balance anomaly detection. In Fig. 1, the red crosses indicate anomalous points detected by the algorithm, specifically flagging the first four standard deviations at $\lambda$=387.85, 391.40, 394.96, and 398.51 nm (blue spectra) as anomalies. The algorithm assists finding out that these four channels actually have zero standard deviations as all their data samples are equal to zero. We observe that this phenomenon is exclusive to L1b radiance and does not occur in the raw data. This suggests that it is a result of the radiometric calibration procedures applied in [31]. Additionally, Fig. 1 reveals that across the dataset, there is a decline in standard deviation values near $\lambda$=760 nm (NIR), primarily due to a sharp reduction in captured light intensity around this wavelength. This decline in light intensity is consistently observed across the image pixels in the dataset. The phenomenon is often associated with atmospheric factors, such as the significant absorption of light by oxygen at this wavelength. The IF algorithm also identifies deviations around $\lambda$=760 nm as anomalous. In particular, it identifies as anomalies the deviations at $\lambda$=734.03, 744.34, 747.77, 758.06, 761.49, and 768.34 nm (channels 99, 102, 103, 106, 107, and 109, starting from 0). Some of these deviations are within the descent peak around $\lambda$=760 nm, while others are just outside, likely due to the 0.08 sensitivity threshold mentioned earlier. In any case, we conclude that channels around the decline at $\lambda$=760 nm may not provide valuable information for pixel-level segmentation. Furthermore, we also add channel 108 ($\lambda$=764.91 nm), as it falls within the descent peak but its deviation remains undetected by the IF algorithm. Consequently, we consider that only the following NIR channels potentially lack meaningful information for segmentation: $\lambda$=758.06, 761.49, 764.91, and 768.34 nm (channels 106, 107, 108, and 109).
Figure 1: Spectral deviations across channels (L1b radiance).
In summary, we suggest removing eight channels from the sea-land-cloud dataset (L1b radiance): the first four with data samples equal to 0 at $\lambda$=387.85, 391.40, 394.96, and 398.51 nm (blue) and the remaining four with a decline at $\lambda$=758.06, 761.49, 764.91, and 768.34 nm (NIR). This reduces the dataset from its original 120 channels to 112 channels, which we utilize in training our ML and DL models in the subsequent methodology. We opt not to exclude further additional channels, as we exercise prudence in our selection to mitigate the risk of inadvertently excluding channels that could have relevance for segmentation. We acknowledge that other methods exist for more accurately identifying channels with valuable information [33, 32], but this falls outside the scope of our current work, which is primarily focused on on-board segmentation.
Since this work additionally studies the impact on the segmentation outcomes when further reducing the number of channels [33], for the particular context of sea-land-cloud HSI segmentation, we utilize the 112-channel data to generate another dataset with significantly fewer channels. Considering that the legacy algorithms for cloud detection in HS satellite imagery typically select 3 channels to mimic RGB image processing [10, 11, 12, 13], we create a new dataset from the original 112 channels, reducing it to merely 3 channels. To achieve this reduction from 112 to 3 channels, we explore the extensive literature on HSI dimensionality reduction techniques, which includes methods like Principal Component Analysis (PCA) [32] and Independent Component Analysis (ICA) [34]. Upon the normalization of the dataset and in alignment with these approaches, we employ PCA to reduce the 112 channels to 3, similarly as in RGB. However, to alleviate to a certain extent the computational demands of satellites operating with tight power constraints, we propose to conduct both the training and inference of PCA exclusively on the ground. Eliminating PCA inference on the satellite enables potential power savings, which could be essential to conduct the subsequent segmentation inference during flight. Consequently, and as also outlined by the literature [11], we suggest implementing on board a direct channel selection (3 bands) instead of deploying PCA inference. In this work, the purpose of employing PCA on ground is not to obtain a low-dimensional dataset directly with the PCA technique, but rather to serve as a tool for identifying the channels containing the most valuable information. In particular, we identify through PCA, for a single principal component, the channels that yield the most substantial influence on the component during the image reduction process. The influence of each channel in the principal component is measured via the component weights (one loading per channel), indicating the particular extent of each channel’s contribution to the component. In the case of the sea-land-cloud dataset (112 channels in L1b radiance), we find that the majority of influential channels are predominantly within the NIR range. However, given the necessity of considering other spectral ranges in addition to NIR for the adequate segmentation of sea, land, and clouds – as the NIR range alone is likely to be insufficient – we opt to pick the top three influential channels (with highest weight loadings) from only the visible blue, second the visible green and red combined, and finally the NIR. In the dataset, the most influential channels within each spectral range are: 412.72 nm (channel 7) for the blue spectrum, 699.61 nm (channel 89) in the green and red spectra, and 747.77 nm (channel 103) in the NIR spectrum. These align closely with the L1b radiance bands frequently analyzed in HYPSO-1, especially for channels 89 and 103.
2.3 Semantic Segmentation: Machine Learning and Deep Learning Models
2.3.1 Processing of 1D vs 3D Data
Our method involves training 16 distinct ML and DL models that are considered to be relevant for on-board processing, with a particular emphasis on DL models, as they are often better suited to capture complex data relationships. To train the models, we maintain the same data split used in our prior article [31], allocating 30 HS images for training (79%), 3 images for validation (8%), and 5 images for testing (13%). Furthermore, we employ an additional deployment set of 10 HS images, also from the sea-land-cloud dataset, to test the best-performing model’s ability to generalize to further new data. Additional details such as the specific image identifiers for the split are provided in the supplementary material referenced at the end of the paper.
The ML and DL models presented in this work handle the data in either one-dimensional (1D) or three-dimensional (3D) formats. On the one hand, when we refer to 1D processing, it implies that the models make predictions solely based on spectral characteristics, disregarding any nearby spatial context. Examples of such models include 1D-CNNs based on 1D convolutions, where the input consists of a spectral signature without any spatial context. Henceforth, within this paper, we refer to these types of models as 1D models. When we refer to 3D processing, it signifies that the models concurrently analyze multiple adjacent spatial pixels, incorporating the hyperspectral datacube with both spectral (1D) and spatial (2D) information to predict the class for each pixel. Models like 2D-CNNs, that we refer henceforth as 2D models serve as examples, where the input is a 3D image containing both spatial and spectral data. Convolutional layers in 2D-CNN models operate by first applying 2D convolutions individually between each spectral channel and its corresponding kernel component. Subsequently, they aggregate the 2D convolutional results stacked in the spectral dimension to produce a final 2D feature map for the neural kernel.
Before introducing the models in the next section, we first describe the necessary arrangement of the data dimensions required for training them. On the one hand, in the case of 1D-CNN (or 1D ML) models, we transform each HS image into a 2D array resembling tabular data. Therefore, these models take inputs formatted as BATCH SIZE $\times$ CHANNELS, where BATCH SIZE indicates the number of image pixels for which predictions are to be inferred. 1D-CNN models perform 1D convolutions separately on each image pixel across its channels. On the other hand, for 2D-CNNs, optimizing memory and computational resources during training and inference is even more essential, and to help achieving this, we divide each HS image datacube into smaller cubes, often referred as 3D patches. First, to guarantee integer spatial dimensions for the patches, we explore both cropping and padding techniques prior to training. We eventually choose padding to retain the integrity of edge information in the images. Padding extends the image edges, which can be conveniently eliminated by cropping the padded spatial dimensions in the segmented image. After image padding, we proceed with non-overlapping patching. Unlike overlapping patching, this approach guarantees that each pixel is processed only once to reduce computational complexity. The input format for the 2D-CNNs is BATCH SIZE $\times$ PATCH SIZE $\times$ PATCH SIZE $\times$ CHANNELS, where BATCH SIZE refers in this case to the number of 3D patches. We find that a PATCH SIZE of 48 is suitable for our experiments, and this decision aligns with our design of the encoder-decoder networks, which consistently decrease the spatial dimensions by a factor of two via pooling. Consequently, selecting a PATCH SIZE of 48 ensures that the spatial dimensions subjected to pooling remain integer as the network’s depth increases.
2.3.2 Baseline Segmentation Models
As a starting point for on-board segmentation, we begin using simple baseline models consisting of traditional ML algorithms, often less complex and more explainable than deep neural networks. For convenience, during this initial phase, the models process 1D data (spectral signatures), which is a simpler processing strategy compared to handling 3D image patches. The initial baseline ML models consist of Stochastic Gradient Descent (referred henceforth as 1D-ML/SGD), Gaussian Naïve Bayes (1D-ML/NB), Linear Discriminant Analysis (1D-ML/LDA), and Quadratic Discriminant Analysis (1D-ML/QDA). Next, we incorporate multiple state-of-the-art DL models for 3D patch processing, such as NU-Net [35] and NU-Net-mod [12], previously employed in satellite imagery for cloud detection in RGB (FACSAT-1) and MSI (Sentinel-2B). Furthermore, we incorporate additional DL models, namely C-UNet and C-UNet++ [36], distinguished by the authors [36] as lightweight for their minimal parameter count and low computational requirements, making them well-suited for deployment on satellites. C-UNet and C-UNet++ are compact versions of the standard encoder-decoder U-Net architecture [37]. The subsequent literature [38] building upon [36], explores C-UNet and C-UNet++ models suggesting their possible use for segmentation in HS missions such as HYPSO-1. However, the method in [38] involves training on only two limited HS images, the conventional Pavia Centre and AeroRIT scene, which use sensors different from HYPSO-1 and do not include any cloud or oceanic data, which is of particular interest for ocean observation in HYPSO-1. In our work, we refer to the compact models in [38] as 2D-CUNet and 2D-CUNet++, and to their further reduced versions in [38] as 2D-CUNet Reduced and 2D-CUNet++ Reduced. Moreover, we add models from the FAUBAI project [39]. FAUBAI, a collaborative project between research and industry for ESA, focuses on segmentation for HS imagery for its possible future use in satellites like HYPSO-1, among other space missions. We refer to the networks in [39] as 2D-UNet FAUBAI and to its reduced version as 2D-UNet FAUBAI Reduced. Although Netteland [39] recommends any of these two networks for tasks such as cloud segmentation in HS satellite imagery, we note that similarly to [38], Netteland [39] conducted model testing only on a restricted set of four conventional HS images, namely, Pavia Centre, Pavia University, Salinas, and Indian Pines, which do not contain cloud or oceanic data.
2.3.3 Customised Segmentation Models: Lightweight CNNs for Optimal On-Board Deployment
2.3.3.1 Review of 1D-CNNs: Liuetal, Huetal, and LucasCNN Models
The Liuetal [40], Huetal [41] and LucasCNN [42] 1D-CNN models have been used before for data analysis in soil spectroscopy to segment various soil texture classes in HS data. We find these models of interest also for satellite-based inference due to their simple 1D processing approach. Indeed, in our previous article [31], we verified that a modified version of Liuetal achieved high performance, surpassing the state-of-the-art in soil spectroscopy [42].
In Table I, we show the parameter counts (assuming 112 channels and 3 classes to segment) for the Liuetal, Huetal and LucasCNN models, highlighting that Liuetal has the lowest number of parameters among the three, although they still have parameter counts in the order of tens of thousands. Therefore, following our pursuit of further reducing the parameter count to merely a few thousands, more convenient for inference aboard satellites, we modify the soil models in [42], while aiming to preserve their performance, in the next sections.
Table I: Overview of hyperparameters and parameter count for 1D-CNN segmentation models.
Model Parameter Count Layers Parameters
Liuetal 22,755 Hidden: 4 1D Convolutions & 4 Max Poolings Output: Flatten & 1 Dense Hidden/Convolutions: N=32,32,64,64; K=3; Act.=ReLU Hidden/Poolings: ÷2 Output/Dense: N=3; Act.=Softmax
1D-Justo-LiuNet 4,563 Hidden: 4 1D Convolutions & 4 Max Poolings Output: Flatten & 1 Dense Hidden/Convolutions: N=6,12,18,24; K=6; Act.=ReLU Hidden/Poolings: ÷2 Output/Dense: N=3; Act.=Softmax
Huetal 100,663 Hidden: 1 1D Convolution & 1 Max Pooling Output: Flatten & 2 Dense Hidden/Convolution: N=20; K=12; Act.=Tanh Hidden/Pooling: ÷2 Output/Dense 1: N=100; Act.=Tanh Output/Dense 2: N=3; Act.=Softmax
1D-Justo-HuNet 9,543 Hidden: 1 1D Convolution & 1 Max Pooling Output: Flatten & 2 Dense Hidden/Convolution: N=6; K=9; Act.=Tanh Hidden/Pooling: ÷2 Output/Dense 1: N=30; Act.=ReLU Output/Dense 2: N=3; Act.=Softmax
LucasCNN 80,155 Hidden: 4 1D Convolutions & 4 Max Poolings Output: Flatten & 3 Dense Hidden/Convolutions: N=32,32,64,64; K=3; Act.=ReLU Hidden/Poolings: ÷2 Output/Dense 1: N=120; Act.=ReLU Output/Dense 2: N=160; Act.=ReLU Output/Dense 3: N=3; Act.=Softmax
1D-Justo-LucasCNN 25,323 Hidden: 1 1D Convolution & 1 Max Pooling Output: Flatten & 3 Dense Hidden/Convolution: N=16; K=9; Act.=Tanh Hidden/Pooling: ÷2 Output/Dense 1: N=30; Act.=Tanh Output/Dense 2: N=5; Act.=Tanh Output/Dense 3: N=3; Act.=Softmax
2D-Justo-UNet-Simple 7,641 Hidden: 3 2D Convolutions & 2 Max Poolings & 2 Upsamplings Output: 1 2D Convolution * Batch normalization after all 2D Convolutions Hidden/Convolutions: N=6,12,6; K=3x3; Act.=ReLU Hidden/Poolings: ÷2 Hidden/Upsamplings: x2 Output/Convolution: N=3; K=3x3; Act.=Softmax
2.3.3.2 Hyperparameter Fine-Tuning Optimization
To modify the soil models by reducing their trainable parameters while not degrading the models’ performance, we fine-tune their hyperparameters following the next procedure. To begin, in the first column of Table I, we list the state-of-the-art soil models in italic font [42], while our customized models are listed in bold. We also differentiate them by the addition of the term “Justo” to their names. In the subsequent columns of the table, for our models, we utilize bold font to highlight the hyperparameters that we modify compared to the soil models. Taking into account the promising nature of 1D-CNNs as lightweight models for on-board deployment, we focus the hyperparameter optimization exclusively on them, and we omit this optimization for 2D-CNN models. The modifications to hyperparameters, which are elaborated on in the following section, are a direct outcome of the subsequent hyperparameter fine-tuning process.
To optimize hyperparameters of 1D-CNNs, we begin by assigning them intuitive values that can yield acceptable performance. Then, we conduct a search using a hyperparameter grid, defined from the previous values, to identify the optimal combination of hyperparameters. This involves training and evaluating the resulting models with cross-validation and selecting the model with the highest overall accuracy. However, considering the high dimensionality of HSI data, we randomly choose only 2 million spectral signatures (approximately 10% of the training data), since training hundreds of models based on the multiple grid combinations would be exceedingly time-consuming, if the complete training set would have been used. Additionally, while evaluating the grid models through cross-validation, we restrict the evaluation to just two fold permutations to further decrease computation time. We note that it is not fully necessary to use the entire training set for hyperparameter tuning during grid search, as our goal is to obtain the optimal hyperparameters, yet within a reasonable time frame. Upon deciding the optimal hyperparameters, during the model training phase presented later, we certainly utilize all training pixels.
2.3.3.3 Introducing New Lightweigth 1D-CNNs: 1D-Justo-LiuNet, 1D-Justo-HuNet and 1D-Justo-LucasCNN Models
Figure 2: Lightweight 1D-Justo-LiuNet architecture for on-board semantic segmentation.
The hyperparameter optimization in previous section aims to achieve optimal accuracy, but we emphasize that the minimization of parameter counts is still the primary focus of our method. In this context, in our prior article [31], we presented a modified version of the Liuetal network resulting in 124,163 parameters, which might be considered large for in-flight deployment. However, our results demonstrated superior performance when contrasted with the Liuetal model [42]. We consider that the large parameter count in [31], which did not result in model overfitting, likely contributed to the significant performance improvement over the state-of-the-art. Nevertheless, in this work, our goal is to once again outperform the state-of-the-art, while also reducing the parameter count, in contrast to our previous work [31]. We aim to achieve a parameter count even lower than the count of the Liuetal model. Therefore, we next present the 1D-Justo-LiuNet (Fig. 2), which achieves 80% reduction in parameters compared to Liuetal (22,755) [42]. 1D-Justo-LiuNet is adapted to only 4,563 parameters, as shown in Table I. This reduction is accomplished by first following the approach in our prior article [31], suggesting to start with a base seed of neural kernels (32 in [31]) and progressively increase the kernel counts (denoted as N in Table I) based on the seed multiplied by factors of 2, 3, and 4 as the network deepens. In this work, we however start with a smaller seed of 6 kernels, instead of 32, aiming to achieve the substantial 80% reduction in the model’s parameter count. Moreover, as we did in our previous publication, we double the kernel size (denoted as K in Table I) compared to [42] in order to compensate for the reduction in the kernel counts, thereby ensuring that performance is not adversely affected.
For the Huetal model [42], the high parameter count in Table I is mainly a result of a fully connected layer with numerous neurons preceding the output classification layer [41]. We refine the model, naming it 1D-Justo-HuNet, by reducing 70% of the neural units in this dense layer. By implementing additional modifications such as reducing the number of kernels and adjusting the kernel size, we attain a significant 90.50% reduction in parameter count (9,543). As for LucasCNN [42], our model, 1D-Justo-LucasCNN, follows the same approach applied in 1D-Justo-HuNet to reduce the parameter count, i.e., we significantly reduce the neural units in the two dense layers before the classification layer. Specifically, LucasCNN has 280 neural dense units, while 1D-Justo-LucasCNN reduces the units by 87% to just 35 neurons. Furthermore, 1D-Justo-LucasCNN includes other adjustments, such as transforming the 4-scale LucasCNN into a single-scale network, significantly reducing the number of kernels while tripling the kernel size, and employing a hyperbolic neural activation instead of a linear function. Table I illustrates a 68.41% reduction in parameter count (25,323) achieved by 1D-Justo-LucasCNN. The architectural details (as in Fig. 2) of the 1D-Justo-HuNet and 1D-Justo-LucasCNN are available in the supplementary materials, which also include the codes to reproduce the models.
2.3.3.4 Introducing a New Lightweight 2D-CNN: 2D-Justo-UNet-Simple Model
Finally, to achieve lightweight models with reduced parameters, not only for 1D-CNN models as in previous section, but also for 2D-CNN architectures, we introduce: 2D-Justo-UNet-Simple (7,641 parameters in Table I). The 2D-Justo-UNet-Simple consists of a compressed version of the encoder-decoder U-Net structure to only 2 scales (as opposed to the typical 4 scales in U-Nets [37]), with a bottleneck level of only 12 kernels. Following each 2D convolution operation and before pooling and upsampling, we insert batch normalization before the rectified linear activation to enhance model generalization and mitigate potential gradient vanishing or exploding effects during training. We note that in the decoder’s contracting path, we employ upsampling as an alternative to utilizing operations like transposed convolutions. While transposed convolutions may enhance performance, they also lead to an increase in the number of trainable parameters. The code and architectural details (as in Fig. 2) of the 2D-Justo-UNet-Simple are provided in the supplementary materials.
2.3.4 Training Procedure
The DL models (implemented in Keras) are trained using an NVIDIA RTX A4000 GPU to increase the training speed, while the ML models (implemented in scikit-learn) are trained on a 13th Gen Intel(R) Core (TM) i9-13900K CPU, which results in a slower training process relying on CPU computation. We employ supervised learning to train the ML and DL models, with a particular focus on the training process for DL models. This is crucial because DL models are more susceptible to overfitting than ML models. We start by using the 112-channel dataset (from Section 2.2) to train the 1D and 2D models. However, when working with the 3-channel dataset, we exclusively focus on training 2D-CNN models. This decision is primarily based on the fact that 1D models do not take spatial information into account and rely solely on the spectral features, which, in this case, comprise only 3 channels. Furthermore, 1D-CNN models can face challenges related to the diminishing channel features as the network depth increases. When working with three channels, the reduction in features is so pronounced that, in the initial layers of the network, the processed features already diminish to zero, rendering prediction for three channels unfeasible.
In our training procedure, we initially select high-level parameters like the number of epochs and the batch size, followed by the choice of the lower-level hyperparameters mentioned in previous sections. For both 1D-CNN and 2D-CNN models, we find that our dataset benefits from small batch sizes, increasing the frequency of parameter updates, but also extending the training time. In addition, we find that it is advisable to utilize a small number of epochs to prevent overfitting. Specifically, for 1D-CNNs, we use 2 epochs and a batch size of 32 pixels, while for 2D-CNNs, we opt for 3 epochs and a batch size of 4 patches. During training, we employ the Adam optimizer (Keras) with its default parameters to minimize the categorical cross-entropy loss, while using training accuracy as evaluation metric. As regards to batch size during inference, Keras dynamically selects the batch size for optimal memory efficiency.
To evaluate the model’s performance at the end of each training epoch, we focus on the overall validation accuracy (validation set imbalance: 51.25%, 28.13%, and 20.62% for sea, land, and clouds, respectively), and we aim to achieve an empirical validation accuracy of around 0.90 while training. After completing the training process, we evaluate false positives and false negatives (inversely proportional) by treating the segmentation of each class as a binary classification problem. We note that the significance of these false positives and false negatives varies by class, as elaborated further in the upcoming section. However, we emphasize in advance that we aim for a reasonably balanced trade-off between false positives and false negatives across sea, land, and cloud classes before selecting the final model. In other words, we attempt to avoid situations where, for instance, we have very low false positives but extremely high false negatives. Therefore, we consider false positives and false negatives intuitively during model training, but we acknowledge more rigorous methods that focus on minimizing either false positives or false negatives by adjusting decision thresholds. These adjustments can involve, for instance, altering the decision boundary during the categorical class assignment based on different probability thresholds, consequently impacting the false positives and false negatives for each class. Additionally, applying penalties within the loss function to address specific misclassifications can also affect false positives and false negatives. However, in our work, we consider sufficient to predict the categorical classes during inference by selecting the class with the highest probability, while equally penalizing all classes during training, since the dataset imbalance is moderate. In the supplementary material, we provide the trained model files with the learned parameters in this work.
Finally, while quantization does contribute to further compacting lightweight models [11], we emphasize that we do not perform parameter quantization. The literature [14] demonstrates that in certain cases, where highly compressed light models are utilized, quantization is unnecessary, as the reduction in parameter count alone may be sufficient for in-orbit deployment.
2.4 Metrics for Segmentation Performance and Example Application: An On-Board Sea-Land-Cloud Ranking System
Figure 3: An intelligent on-board processing pipeline integrating a ranking system based on sea-land-cloud segmentation. $T\!H_{cl}$, $T\!H_{sea}$, and $T\!H_{land}$ represent thresholds configured at the operations center to regulate data management during in-flight operations.
To assess the models’ performance after training, we employ various metrics to evaluate the quality of the segmented results on the test set. First, our primary focus is on average accuracy, a widely used metric for evaluating segmentation models. Unlike overall accuracy, this metric is particularly convenient for imbalanced test sets, providing a more realistic assessment. Furthermore, adhering to the literature of cloud detection in satellite imagery, we also calculate the false and true alarm ratios. For this, as previously mentioned, we treat the segmentation of each class as a distinct binary classification problem. Finally, in our supplementary material further supporting the conclusions of our work, we present additional state-of-the-art metrics that we consider secondary in importance compared to the previously mentioned metrics. Specifically, these metrics include the F-1 score in addition to the overall accuracy.
Next, following the assessment of our segmentation results, we introduce example applications to study the utility of the sea, land, and cloud segments in increasing the satellite’s autonomy by decreasing postprocessing demands and reducing decision-making latency. In Fig. 3, we propose a novel sea-land-cloud ranking system for satellite imagery. This system prioritizes HS images for downlink based on sea, land, and cloud coverage levels, ensuring high-priority transmission. The images are arranged in ascending order of cloud coverage levels and descending order of sea and land coverage levels, with the first positions reserved for images of higher priority. Unlike existing ranking systems that target only clouds [12], our approach is the first of its kind in the field, simultaneously identifying and ranking water, land, and clouds within priority queues for HSI-equipped satellites. In Fig. 3, we also suggest other use cases that further highlight the usefulness of the sea-land-cloud segments. These use cases include alternatives such as discarding or applying lossy compression to non-essential pixels at the satellite’s edge to enhance the efficient utilization of the Shannon channel capacity, particularly vital for satellites like HYPSO-1 that may be operating with limited line-of-sight (LOS) towards ground stations [26].
We introduce additional metrics for the ranking system to assist the evaluation of the ML and DL models. To assess the predicted rankings for each coverage level, we use the Spearman’s rank correlation coefficient. This measures the relationship between the predicted and ground-truth rankings, with values ranging from 1-1 to 1, representing a perfect match at the upper end. Moreover, we compute the Mean Absolute Error (MAE) for each coverage level, treating each level as a separate regression task. To maintain simplicity, we include the details about MAE in our supplementary materials.
In the context of on-board data management (also in Fig. 3), the false alarm ratios due to misclassifications indicate either the potential loss of data or the transmission of irrelevant data during downlink, thereby diminishing downlink efficiency. In cloud segmentation, the True Positive Ratio (TPR) involves correctly handling cloudy pixels, while the True Negative Ratio (TNR) relates to recognizing non-cloudy data with potential value. The False Positive Ratio (FPR) is crucial because it indicates misclassified pixels, wrongly identified as clouds but that actually correspond to sea or land. This can result in the elimination or lossy compression of valuable data. On the other hand, the False Negative Ratio (FNR) is less critical than FPR, as it does not involve data loss, but only an inefficient usage of the downlink transmitting undetected pixels actually corresponding to clouds. In the context of sea segmentation, true positives signify correct identification of valuable water data, while true negatives correctly recognize non-relevant data for potential deletion or compression. In sea-focused campaigns, FNR is however critical, as it deals with misclassified sea-related pixels, potentially leading to loss of valuable data. In contrast, FPR is of less significance since its impact lies only in downlink efficiency through the transmission of irrelevant pixels (clouds or land), rather than causing actual data loss. The same analysis and conclusion apply when examining the context of land segmentation. In summary, without taking into account the earlier suggested compromise between false positives and false negatives, for both sea and land categories, FNR has priority over FPR due to a likelihood for valuable data loss in the case of high FNR. However, for cloud data, FPR is critical as it leads to irreversible data loss, with FNR affecting merely downlink efficiency.
Table II: Metrics for segmentation models trained using 112 spectral channels.
Model Accuracy Spearman’s Coefficient Parameter Count Inference Time [ms]
1D-ML/SGD 0.91 0.90 - 72
1D-ML/NB 0.71 0.27 - 1,234
1D-ML/LDA 0.89 0.93 - 190
1D-ML/QDA 0.89 0.63 - 2,024
1D-Justo-LiuNet 0.93 1.00 4,563 15,136
1D-Justo-HuNet 0.91 1.00 9,543 10,966
1D-Justo-LucasCNN 0.91 0.90 25,323 11,554
2D-NU-Net-mod 0.8 0.83 32,340 364
2D-NU-Net 0.81 0.77 40,332 374
2D-CUNet 0.72 0.83 67,299 272
2D-CUNet++ 0.83 0.77 24,619 294
2D-CUNet Reduced 0.76 0.90 22,019 270
2D-CUNet ++ Reduced 0.90 0.83 12,379 294
2D-UNet FAUBAI 0.88 0.90 26,534,211 462
2D-UNet FAUBAI Reduced 0.85 0.83 1,956,835 388
2D-Justo-UNet-Simple 0.92 0.93 7,641 318
Table III: Ranking of models’ metrics from best to worst for segmentation models trained over 112 channels.
Ranking Accuracy Spearman’s Coeff. Parameter Count Inference Time
1 1D-Justo-LiuNet 1D-Justo-LiuNet 1D-Justo-LiuNet 1D-ML/SGD
2 2D-Justo-UNet-Simple 1D-Justo-HuNet 2D-Justo-UNet-Simple 1D-ML/LDA
3 1D-ML/SGD 2D-Justo-UNet-Simple 1D-Justo-HuNet 2D-CUNet Reduced
4 1D-Justo-HuNet 1D-ML/LDA 2D-CUNet++ Reduced 2D-CUNet
5 1D-Justo-LucasCNN 1D-Justo-LucasCNN 2D-CUNet Reduced 2D-CUNet++
6 2D-CUNet++ Reduced 1D-ML/SGD 2D-CUNet++ 2D-CUNet++ Reduced
7 1D-ML/LDA 2D-CUNet Reduced 1D-Justo-LucasCNN 2D-Justo-UNet-Simple
8 1D-ML/QDA 2D-UNet FAUBAI 2D-NU-Net-mod 2D-NU-Net-mod
9 2D-UNet FAUBAI 2D-CUNet 2D-NU-Net 2D-NU-Net
10 2D-UNet FAUBAI Reduced 2D-NU-Net-mod 2D-CUNet 2D-UNet FAUBAI Reduced
11 2D-CUNet++ 2D-CUNet++ Reduced 2D-UNet FAUBAI Reduced 2D-UNet FAUBAI
12 2D-NU-Net 2D-UNet FAUBAI Reduced 2D-UNet FAUBAI 1D-ML/NB
13 2D-NU-Net-mod 2D-CUNet++ - 1D-ML/QDA
14 2D-CUNet Reduced 2D-NU-Net - 1D-Justo-HuNet
15 2D-CUNet 1D-ML/QDA - 1D-Justo-LucasCNN
16 1D-ML/NB 1D-ML/NB - 1D-Justo-LiuNet
Table IV: Metrics for segmentation models trained using only 3 spectral channels.
Model Accuracy Spearman’s Coefficient Parameter Count Inference Time [ms]
2D-NU-Net-mod 0.66 0.37 32,231 224
2D-NU-Net 0.80 0.53 32,375 192
2D-CUNet 0.86 0.83 59,451 162
2D-CUNet++ 0.84 0.6 16,771 100
2D-CUNet Reduced 0.87 0.73 14,171 156
2D-CUNet ++ Reduced 0.84 0.83 4,531 122
2D-UNet FAUBAI 0.87 0.77 26,471,427 346
2D-UNet FAUBAI Reduced 0.85 0.57 1,941,139 230
2D-Justo-UNet-Simple 0.85 0.83 1,755 112
Figure 4: Comparison of false cloud detection ratios for segmentation models trained over 112 channels.
Figure 5: Comparison of false sea detection ratios for segmentation models trained over 112 channels.
Figure 6: Comparison of false land detection ratios for segmentation models trained over 112 channels.
3 Results
We conduct model training using 112 channels of both raw data and L1b radiance in the test set (further discussed in Section 4). Table II presents the outcomes of the segmentation models trained on 112 channels in L1b radiance. In the table, we utilize bold font to accentuate the novel models. The table displays the models’ performance for segmentation, with the primary metric being the average accuracy across the sea, land, and cloud classes. Additionally, to assess the sea-land-cloud segmentation for the context of the on-board ranking system, we also include the average Spearman’s rank coefficient as performance metric. In particular, the mean coefficient across the respective rankings based on sea, land, and cloud coverage levels. Moreover, considering the extensive number of parameters in DL models, which in general can be a challenge for on-board implementations, we also include information about the network’s parameter count to offer insights of the potential memory demands for in-orbit deployment. In addition, we provide the average inference time per HS image measured on ground, on the NVIDIA RTX A4000 GPU (for DL), and on the 13th Gen Intel(R) Core (TM) i9-13900K CPU (for ML). From the numbers presented in Table II, we sort the models in Table III based on their respective metrics. In Table III, we first sort the models solely based on their average accuracy, from highest to lowest, and subsequently, we sort them in the following columns, first by the Spearman’s coefficient (highest to lowest), then by the number of parameters (lowest to highest), and finally by inference time (shorter to longer). Therefore, the first rows in the table represent the top models for each respective metric. However, as the average accuracy only provides a broad overview of the mean segmentation performance across the sea, land, and cloud classes, we additionally provide fine-grained insights into the models’ performance for each separate class by presenting their false detection ratios.
In Figs. 4, 5, and 6, we provide a comparative visual representation of false alarm ratios for the segmentation of each distinct class. We note that the top-performing models get positioned in the lower-left corner, indicating the optimal trade-off balance between false positives and false negatives. Within the context of the example case (Section 2.4), the graphs illustrate the false alarm ratio on the Y-axis corresponding to the loss of valuable data, while the X-axis pertains to a false detection ratio that reflects the inefficient utilization of the downlink. In relation to this trade-off, we stress that while minimizing the loss of valuable data remains crucial, we also highlight the importance of maintaining an acceptable downlink efficiency. Finally, our experiments conclude by retraining the models using the 3-channel dataset. We employ this dataset to assess the impact of the channel reduction (Section 2.2) on the models’ performance, parameter count, and inference times, as shown in Table IV. We note that the table omits the 1D models due to the 3-channel dataset’s restricted spectral information, as mentioned earlier (Section 2.3.4). Notably, the empirical results are further confirmed by retraining once again the models with a different set of three spectral channels. These new results align with the previous findings. For simplicity, we include these new figures in the supplementary material. As previously noted, the dataset predominantly contains instances of sea and land, with clouds representing the minority class. Consequently, the test set also exhibits a lower proportion of cloud data (14.13%), in contrast to sea (39.01%) and land (46.85%). We recognize this as a limitation that demands future refinement. To mitigate the impact of the class imbalance during the testing phase, we conduct additional experiments. Namely, we perform inference on an additional deployment set specifically designed to test the model’s generalization on e.g. cloudy images to further support the findings of our work, and we provide the results in the supplementary material.
4 Discussion
4.1 Analysis of Results for Models Trained on 112 Spectral Channels
We first analyze the results for the models trained on 112 channels (Tables II and LABEL:Table:models_sorted_according_to_performance). Our objective is to investigate how the utilization of spatial context by 2D models affects the sea-land-cloud segmentation across test images. We assess the models’ strengths and limitations in terms of performance, spanning from the average accuracy and the false alarm ratios to the Spearman’s correlation coefficient for the ranking system. We finalise this analysis by discussing the model’s parameter count and inference time, which are relevant metrics for on-board deployment. Prior to the analysis of all the metrics, we intuitively formulate a hypothesis, independent of the empirical findings. The hypothesis asserts that it is reasonable to anticipate better performance for 2D models compared to 1D models. This expectation arises because 2D models, unlike 1D models relying solely on spectral context, analyse a richer context comprising both spectral and spatial information. We proceed next to elaborate whether the empirical results substantiate such hypothesis.
4.1.1 Segmentation Performance (I): Average Accuracy
When examining the segmentation performance solely based on the average accuracy across sea, land, and cloud classes, we observe that our proposed model, 1D-Justo-LiuNet, achieves the highest accuracy of 0.93 (Table II). The result of a 1D model outperforming 2D models is not an exception, as four out of five models ranking among the top performers in accuracy in Table LABEL:Table:models_sorted_according_to_performance are consistently 1D models. The only exception in this top five ranking is our 2D model, 2D-Justo-UNet-Simple, which achieves an accuracy of 0.92, closely approaching the performance of the 1D-Justo-LiuNet. These findings lead to the conclusion that the initially expected hypothesis suggesting that 2D models would outperform 1D models is not valid, as the 2D models actually perform worse. This phenomenon can be attributed to the fact that 1D models, unlike 2D using the spatial context, are not affected by the modest spatial resolution of HYPSO-1’s images while not suffering from the spatial variations that may be practically irrelevant for pixel-level segmentation. Instead, 1D models dedicate their efforts to analyzing exclusively the spectral context, which is a valuable domain in HS imagery. Additionally, Table LABEL:Table:models_sorted_according_to_performance demonstrates that our proposed models (1D-Justo-LiuNet, 2D-Justo-UNet-Simple, 1D-Justo-HuNet, and 1D-Justo-LucasCNN) outperform, in average accuracy, the state-of-the-art models under analysis, when trained on the HYPSO-1 dataset for sea-land-cloud segmentation. Indeed, we surpass models such as the 2D-NU-Net-mod, 2D-NU-Net, 2D-UNet FAUBAI, and 2D-UNet FAUBAI Reduced previously employed in the field of cloud segmentation for remote sensing imagery [12, 39]. Finally, we highlight that the average accuracy of 1D-Justo-LiuNet, 1D-Justo-HuNet, and 1D-Justo-LucasCNN (0.93, 0.91, and 0.91, respectively) also surpass the literature’s average accuracy for soil texture. As a matter of fact, according to [42], the state of the art achieves significantly lower average accuracy of 0.59, 0.61, and 0.56 for Liuetal, Huetal, and LucasCNN, respectively.
4.1.2 Segmentation Performance (II): The Effect of L1b Radiance Calibration
The models in Table II are trained on L1b calibrated radiance data. L1b radiance is one of the standard data formats in HS imagery, which not only allows satellite comparisons but also removes sensor-specific characteristics and biases. Next, we empirically examine how L1b radiance calibration impacts segmentation performance. To be brief while supporting the next discussion, we include the average accuracy for the models trained on raw data in the supplementary materials of this article to allow the comparison with the models trained on L1b radiance in Table II. For the highest-performing DL models in terms of average accuracy shown in Table III (1D-CNNs), they exhibit slightly superior performance with L1b radiance calibration compared to training them with raw data. A similar trend occurs for the 2D models ranked just below the 1D models, which also demonstrate enhanced performance with radiance data versus raw data. Conversely, the models at the lower end of the accuracy ranking tend to perform better with raw data than radiance calibration. Concentrating only on the superior models in Table III, which hold more significance as they are the top performers, we conclude that L1b radiance calibration tends to positively impact segmentation performance. We attribute the improvement to the correction of errors and biases during L1b radiance calibration, allowing top-performing models to focus on inherent data patterns rather than overfitting on sensor-induced characteristics, leading to improved generalization performance. Nonetheless, we strongly recommend conducting comparative testing on both calibrated and raw data when training the models on datasets distinct from this paper to verify whether calibrated data, such as radiance or reflectance, continues to enhance model performance.
4.1.3 Segmentation Performance (III): False Detection Ratios, Data Loss, and Downlink Efficiency
For a more in-depth assessment of the segmentation performance for each separate class, we scrutinize which models are positioned in the lower-left corner in the scatter plots in Figs. 4, 5, and 6. The bottom-left region ensures the most optimal false detection ratios, and for the example system in Fig. 3, the region achieves minimal loss of valuable data, while also minimizing downlink inefficiency. Next, we comparatively analyze the top-performing models located in the lower-left corner of Figs. 4-6.
As previously explained, lower data loss (in the Y-axis) takes precedence, while the downlink inefficiency (in the X-axis) is of lower priority since it does not result in any loss of valuable data. However, as discussed in Section 2.4, we also aim for a reasonably balanced relationship between false positives and false negatives to prevent scenarios where a model achieves a convenient minimal data loss, at the expense of extremely degrading the downlink efficiency.
First, we assess the models in terms of data loss. In Fig. 4 for cloud segmentation, models such as the 2D-Justo-UNet-Simple excels in minimizing the loss of data, approaching nearly no data loss (ideal). This signifies that the 2D-Justo-UNet-Simple outperforms by not misclassifying valuable non-cloud data (sea or land) as clouds, even if it means some cloudy data goes undetected, reducing the downlink efficiency. With respect to sea segmentation (Fig. 5), the 2D-UNet-FAUBAI-reduced stands out for its minimal loss of sea pixels (FNR=0.01). This indicates that there is only a 1% proportion of sea pixels that remain undetected, potentially resulting in their loss or degradation. This outperforms other models such as the 1D-Justo-LiuNet with 5% loss instead. Finally, in Fig. 6 for land segmentation, we observe a remarkable similarity in false alarm ratios in the lower-left optimal region, resulting in a slight overlap. Therefore, we next clarify that the optimal models, sorted in ascending order of FNR (data loss) values, are as follows: 0.02 (2D-Justo-UNet-Simple), 0.03 (1D-Justo-LiuNet), 0.04 (1D-Justo-HuNet), and 0.06 (2D-CUNet++ Reduced). Consequently, the 2D-Justo-UNet-Simple is the most optimal at minimizing the loss of land data, achieving only a 2% ratio of undetected land pixels.
Second, to jointly maintain a reasonably acceptable downlink efficiency and low data loss, we study the models considering the trade-off between both loss of data and downlink efficiency objectives. One could consider the use of Pareto fronts, which are typically applied in multi-objective optimization to address trade-offs between inversely proportional variables. Nevertheless, for simplicity, we measure the trade-off using the Euclidean distance metric that gives equal weight to the two objectives. Specifically, we calculate the distance of each model from the origin point, representing the ideal scenario with zero false alarms. First, in cloud segmentation (Fig. 4), the models closest to the ideal performance have (FPR, FNR) values of (0.02, 0.15) for 1D-Justo-LiuNet and (0.00, 0.17) for 2D-Justo-UNet-Simple, resulting in approximate distances of 0.151 and 0.170, respectively. In brief, 1D-Justo-LiuNet exhibits the shortest distance to the ideal model. Second, the top models for sea segmentation (Fig. 5), with (FNR, FPR) of (0.08, 0.01) for 1D-Justo-LucasCNN, (0.05, 0.02) for 1D-Justo-LiuNet, and (0.04, 0.03) for 2D-UNet-FAUBAI, have approximate distances of 0.081, 0.054, and 0.050, respectively. This suggests a marginally better performance of 2D-UNet-FAUBAI, with a difference of 0.004 in distance compared to the 1D-Justo-LiuNet. However, Table II notes that the 2D-UNet-FAUBAI has over 26 million parameters, whereas our network, 1D-Justo-LiuNet, has fewer than 5 thousand parameters, which is more suitable for on-board deployment. Considering the nearly identical performance and the large parameter difference against FAUBAI’s network, 1D-Justo-LiuNet remains the top model for sea segmentation. Finally, for land (Fig. 6), the optimal segmentation models are of (FNR, FPR) = (0.02, 0.06) for 2D-Justo-UNet-Simple, (0.03, 0.04) for 1D-Justo-LiuNet, (0.04, 0.04) for 1D-Justo-HuNet, and (0.06, 0.04) for 2D-CUNet++ Reduced, with approximate distances of 0.063, 0.050, 0.057, and 0.072, respectively. In summary, 1D-Justo-LiuNet outperforms the other models for land segmentation.
Considering the initial analysis focused solely on data loss and the subsequent analysis regarding the trade-off between data loss and downlink efficiency, we next provide our conclusion. The 2D-Justo-UNet-Simple, with approximately 7 thousand parameters, minimizes the loss of valuable data for most classes. The only exception is the sea class, where it incurs a 5% data loss, in contrast to the 1% loss by 2D-UNet-FAUBAI-reduced which uses around 2 million parameters and is less suitable for on-board inference. However, when we equally consider both data loss and efficient downlink usage, our model, 1D-Justo-LiuNet, excels in minimizing data loss, achieving rates of 2%, 5%, and 3%, while demonstrating relatively minimal inefficiency rates of 15%, 2%, and 3% for cloud, sea, and land classes, respectively. These findings suggest that when employing 1D-Justo-LiuNet, efficiency slightly deteriorates due to a less precise detection of clouds, leading to an overabundance of cloud data that would be transmitted from the satellite (15%). We attribute this issue to the dataset’s class imbalance, where the cloud class constitutes the minority. Nevertheless, the over-detection of sea (2%) and land (3%) pixels is smaller, thereby maximizing the efficiency of the downlink. Additionally, in our supplementary material, we note that model misclassifications resulting from false alarms predominantly manifest spatially at class edges, where pixels are predicted with considerably lower confidence. This phenomenon is primarily attributed to the inaccuracy of the semi-automatic pixel-level annotation process at class boundaries outlined in our previous publication [31].
Before concluding the evaluation of the models’ segmentation performance, we conduct a final analysis of the ranking system to evaluate its effectiveness in sorting images, as explained next. We center our analysis on the Spearman’s coefficient (Table II) to determine which model precisely ranks the HS images based on the priority levels of sea, land, and cloud coverage calculated from the segmented images. According to Tables II and LABEL:Table:models_sorted_according_to_performance, the experimental results confirm that 1D-Justo-LiuNet and 1D-Justo-HuNet are the top performers, surpassing again the state-of-the-art models. As a matter of fact, Table II shows that 1D-Justo-LiuNet and 1D-Justo-HuNet are the only models capable of reaching the ideal Spearman’s coefficient of 1.00, indicating the perfect sorting of the HS images in the rankings, without any errors compared to the ground-truth ranks for sea, land, and cloud levels. This emphasizes the selection of 1D-Justo-LiuNet (or possibly 1D-Justo-HuNet) as the preferred choice for the on-board ranking system.
Furthermore, the results reveal that our 2D model, 2D-Justo-UNet-Simple, although not achieving the same level of ranking performance as the 1D models, also surpass most of their fellow 2D models from the state of the art. Nonetheless, our model, 2D-Justo-LucasCNN, does not produce improved results, as it only yields the same figures to those achieved by other state-of-the-art models (1D-ML/SGD, 2D-CUNet Reduced, and 2D-UNet FAUBAI).
4.1.4 Segmentation Performance (IV): Summary
In light of the previous sections, in addition to the key finding that 1D models outperform their 2D counterparts, we conclude that our 1D-Justo-LiuNet model consistently outperforms the current state-of-the-art models, all trained for the same sea-land-cloud segmentation task in HYPSO-1 data. Specifically, the 1D-Justo-LiuNet model surpasses other models in the literature in terms of segmentation performance, assessed by average accuracy and false detection ratios for each class. Furthermore, these superior results for the 1D-Justo-LiuNet are consistent even in the context of the ranking system application case, and also when compared with the soil texture literature [42]. However, we stress that our results and conclusions are specific to the sea-land-cloud dataset [31], and we acknowledge that model performance may differ when the data source is altered in any manner. For instance, a model that performs well on one dataset may not excel on another. Consequently, to facilitate testing of the models used in this work on other data sources, we provide the software code for all models in the supplementary materials accompanying this paper.
4.1.5 Parameter Count
Following the analysis of the models’ performance, we next examine the number of model parameters that need to be uplinked to a satellite for in-orbit inference (Table II). Our lightweight model 1D-Justo-LiuNet surpasses the existing literature in terms of parameter count (Table LABEL:Table:models_sorted_according_to_performance), ranking at the top, as the model with the fewest parameters, even when compared to other lightweight models in the literature, such as 2D-CUNet, 2D-CUNet++, 2D-CUNet Reduced, and 2D-CUNet++ Reduced [38]. Hence, the lightweight 1D-Justo-LiuNet model stands out as the most practical choice when solely considering the reduction in on-board memory required for storing model parameters. Moreover, despite some of our other models, such as 1D-Justo-HuNet and 2D-Justo-UNet-Simple, exhibit a slight increase in the number of parameters compared to 1D-Justo-LiuNet, they still maintain a parameter count below ten thousands, which is still lower than the parameters of the remaining state-of-the-art models. As mentioned earlier, the lowest-ranked 2D-UNet-FAUBAI network utilizes over 26 million parameters. Even its reduced version, 2D-UNet-FAUBAI-reduced, has nearly 2 million parameters, which we find excessive for on-board deployment while it does not improve the model’s performance neither in terms of segmentation accuracy nor in terms of image ranking. Finally, we emphasize that a significant majority of our top models in terms of performance (accuracy and Spearman’s coefficient in Table LABEL:Table:models_sorted_according_to_performance), occupy the best ranks in terms of parameter count. Therefore, we can assert that we achieve the objective outlined in Section 2 regarding surpassing the state-of-the-art performance, while significantly compressing the models.
4.1.6 Inference Time
The final metric of our analysis is the average inference time per HS image (Table II). Despite the strengths of the 1D models, with particular emphasis on the favorable results by the 1D-Justo-LiuNet, 1D models exhibit a drawback of extended inference times on the processing architecture used in the tests, even considering their lowest parameter count. In Table II, we observe that both 1D ML and DL models, especially DL models with increased intrinsic complexity, exhibit lengthier inference times. In fact, the results reveal that the 1D-Justo-LiuNet has the lengthiest inference time, at around 15 seconds for a single HS image. The results demonstrate, however, that 2D models significantly decrease the inference time, reducing it to the order of milliseconds. The disparity in speed may be attributed to the GPU being primarily optimized for high parallelization of 3D imaging data, which is a prevalent format in computer vision. Consequently, GPUs may exhibit greater efficiency when processing 3D image patches as opposed to 1D sequential data (spectral signatures).
Although satellites like HYPSO-1 experience data latency lasting minutes to hours [26], we propose the following solution if inference times in the order of seconds (e.g., 1D-Justo-LiuNet) are perceived as lengthy for in-orbit inference. Employing the HYPSO-1 satellite as an example case, its On-board Processing Unit (OPU) integrates an FPGA System-on-Chip (Zynq-7030/XC7Z030-1) [9]. This system features the ARM Cortex-A9 processor and provides superior FPGA resources in comparison to other platforms, such as the Zynq-7020, previously employed in cloud detection within satellite imagery [13]. As future work, we suggest implementing the 1D-Justo-LiuNet in software on the processor side to confirm whether the inference time is still lengthy. We anticipate that the processor’s lower-level implementation may yield shorter inference times compared to running inference in higher-level languages like Python. If the inference time would still remain too long, we recommend a combined software-hardware co-design approach. This approach accelerates computationally intensive tasks for faster inference, while maintaining low power consumption using the FPGA fabric logic.
4.2 Analysis of Results for Models Trained on 3 Spectral Channels
We finalize the discussion section by addressing the impact of reducing the number of channels from 112 to 3 on the reported metrics. As indicated in Table IV, when we retrain the models using the 3-channel dataset instead, the inference time decreases compared to the models trained with 112 channels, as there is now less data to process. Moreover, there is a reduction in the number of model parameters due to the smaller number of components within each neural kernel used for computing the corresponding 2D convolutions across the spectral dimension. Nevertheless, the reduction in feature channels leads to a decreased segmentation performance, with the highest accuracy reaching 0.87, unlike the 0.93 top accuracy for the models trained with 112 channels. This decline in performance is also evident in the evaluation of the ranking system, for which none of the models achieve the ideal coefficient of 1.00, as observed earlier with 112 channels. As a matter of fact, the highest Spearman’s coefficient is notably lower at 0.83. Although the main emphasis of our analysis is to examine the impact of reducing the channels from 112 to 3, we shortly discuss next how the 2D models compare to each other for solely 3 channels instead. Table IV shows that the 2D-Justo-Unet-Simple model achieves an accuracy of 0.85, which is close to the highest accuracy of 0.87. In the ranking system, our model exhibits the same top Spearman’s coefficient as 2D-CUNet and 2D-CUNet++ Reduced. Remarkably, the 2D-Justo-UNet-Simple outperforms both 2D-CUNet and 2D-CUNet++ Reduced in terms of both parameter count and inference time.
5 Conclusion
In this work, we conducted experiments including the training and comparison of 16 ML and DL models for on-board deployment using the Sea-Land-Cloud-Labeled dataset from the HYPSO-1 mission [31], as an illustrative case. We demonstrated that for segmenting sea, land, and cloud categories in hyperspectral satellite imagery, 1D models surpass 2D models in terms of performance. This observation also holds for our proposed ranking system combining decision-making for sea, land and cloud coverage. Our model 1D-Justo-LiuNet consistently delivers better results than the state-of-the-art models. Furthermore, we also found that using L1b calibrated radiance data, instead of raw captured data, improved segmentation performance. Moreover, training models with only three spectral channels led to shorter inference times and reduced parameter counts, at the cost of lower segmentation performance. Finally, as future work, we suggest conducting experimental tests to deploy in orbit the 1D-Justo-LiuNet model.
Declaration of Competing Interest
The authors declare that they have no known competing interests that could have appeared to influence the work reported in this paper.
Data availability
The labeled dataset to reproduce the findings in this paper is openly available to download at https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset/. Moreover, we provide as supplementary materials: software codes for the models, trained model files with the learned parameters, information on training-validation-testing, additional evaluation metrics, and generalization tests. The material can be found in our GitHub repository at https://github.com/jonalvjusto/s_l_c_segm_hyp_img.
Acknowledgements
The research leading to these results has received funding from the NO Grants 2014 – 2021, under Project ELO-Hyp contract no. 24/2020, and the Research Council of Norway grant no. 223254 (AMOS), no. 328724 (Green Platform) and no. 325961 (HYPSCI). We extend our gratitude to Anum Masood, Per G. Kjeldsberg, and Simen Netteland.